OpenStudy (anonymous):

help solve y^2+14=10y

OpenStudy (anonymous):

y^2-10y+14=0 use quadratic formula ay^2+by+c=0 where a=1 b=-10 c=14 quadratic formula is \[\Large y=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (mathstudent55):

Subtract the 10y term over to the left side, then use the quadratic formula.

OpenStudy (whpalmer4):

Or if you prefer, complete the square: Put all the terms with powers of y on the left, and numbers on the right \[y^2-10y = -14\]Now take half the coefficient of the y term, square it, and add to both sides \[y^2 -10y + (\frac{-10}{2})^2 = -14 + (\frac{-10}{2})^2\]Rewrite the left side as a square \[(y-5)^2 = 11\]Take square root of both sides\[y-5 = \pm \sqrt{11}\]Solve for y\[y=5\pm\sqrt{11}\]

OpenStudy (whpalmer4):

Important: to complete the square in this manner, you *must* have the coefficient of the squared term = 1. If it isn't, divide the equation by the coefficient.